OpenStudy (anonymous):

suppose that tan t = 5/12 and sin t<0. Find sin (2t) by hand using identities only

OpenStudy (adamaero):

You should memorize this: sin2x = 2sinxcosx

OpenStudy (anonymous):

I know that equation, but I do not know how to use it with this equation.

OpenStudy (adamaero):

@johnweldon1993 @wio @Zarkon

OpenStudy (anonymous):

\[ \tan t= \frac 57 \implies \frac{\sin t}{\cos t} = \frac 57 \]

OpenStudy (anonymous):



OpenStudy (anonymous):



OpenStudy (anonymous):

\[ \sin t = \frac 5h,\ \cos t = \frac 7h \]